# Calibrating dimension reduction hyperparameters in the presence of noise

### v3

@@ 4.2 Practical Examples
In addition to simulated datasets, we looked at three practical datasets: a cytometry by time-of-flight (CyTOF) dataset [15], a single-cell RNA sequencing dataset [16], and a microbiome dataset [2]. For each dataset, we compared the optimal perplexity for locally replicating the original data versus just the signal at varying numbers of signal dimensions. We explore the CyTOF dataset in detail here. Summaries for the other two practical examples can be found in Table 1, and the data processing details for the other two practical examples can be found in the Supplementary Information.
The CyTOF dataset contained 239,933 observations in 49 dimensions [15]. To reduce the computational load, a subset of 5,000 observations was sampled. In line with the ordinary t-SNE workflow, a log transformation was followed by a PCA pre-processing step to reduce the number of dimensions to 30, which still retained 77% of the variance in the original data. The processed data set to be studied consisted of 5,000 observations in 30 dimensions,
$Z+\epsilon\in\mathbb{R}^{5,000\times 30}.$
To determine the dimension of the signal, we drew a scree plot (Figure 5), which depicts flatening after the fifth eigenvalue. And thus, $Y$ was extracted by taking the first five principal components,
$Y=\textrm{PCA}_{5}(Z+\epsilon).$
We computed the t-SNE representations for perplexity values ranging from 10 to 300. For each perplexity value, 20 different t-SNE representations were computed. Figure 6a contains the plots for $f_{\textrm{Shep}}(Z+\epsilon,X)$ vs. $f_{\textrm{trust}}(Z+\epsilon,X)$ and $f_{\textrm{Shep}}(Y,X)$ vs. $f_{\textrm{trust}}(Y,X)$.
@@
As with the simulated examples, there is a difference in trend when the frame of reference is the original data versus when it’s just the underlying signal (Figure 6b). When compared against the original data, trustworthiness is maximized at a perplexity of 50, which is consistent with [10]’s recommendation of setting perplexity to $n/100$. When compared against the underlying signal, trustworthiness is maximized at a larger perplexity of 220, reinforcing the hypothesis that lower values of perplexity may be overfitting the noise.
Figure 7: Trustworthiness vs. Perplexity for $r=8$ (CyTOF)
@@
If we, instead, decide to be a little more conservative and use the first eight principal components to represent the signal, we still see a similar trend (Figure 7). Trustworthiness still increases then decreases with perplexity. When compared against the original data, trustworthiness is maximized at a perplexity of 45. When compared against the signal, trustworthiness is maximized at a perplexity of 65. By including three extra principal components in the signal, we’re assuming the data contains less noise, allowing the model to be more aggressive during the fitting process.
@@ 4.4 UMAP and n_neighbors
If n_neighbors functions similarly to perplexity, we’d expect small values of n_neighbors to overfit the data as well. An identical experiment was run using the Python package umap-learn on the CyTOF data. n_neighbor values ranging from 10 to 300 were tested on the same CyTOF data set. An n_neighbors value of 120 maximized trustworthiness when comparing against the original data, but an n_neighbors value of 160 maximized trustworthiness when comparing against the underlying signal. See Supplementary Information for plots.
@@ 5 Application
With a signal constructed, it remains to compute t-SNE outputs at varying perplexities. It’s recommended that at least a couple outputs are computed for each perplexity to account for t-SNE’s inherit randomness. For each output, one must calculate the trustworthiness and Shepard goodness with respect to the signal. From there, one can choose the representation with the desirable balance of local and global performance. A summary is given in Algorithm 1. Sample code is available at https://github.com/JustinMLin/DR-Framework/.
Optimal Perplexity
Dataset $n$ $p$ $r$ signal + noise signal
Links [20] 500 10 3 30 110
Trefoil [20] 500 10 3 30 110
Mammoth [19] 500 10 3 20 100
High-Dimensional Clusters 210 60 10 40 50
CyTOF [15] 5,000 30 5 50 220
CyTOF [15] 5,000 30 8 45 65
scRNA-seq [16] 864 500 5 40 120
scRNA-seq [16] 864 500 10 50 60
Microbiome [2] 280 66 5 50 90
Microbiome [2] 280 66 8 60 85
Table 1: Summary of results
It is worth noting that computational barriers may arise, especially for very large data sets. To alleviate such issues, trustworthiness and Shepard goodness can be approximated by subsampling before calculation. Furthermore, t-SNE is generally robust to small changes in perplexity [17], so checking a handful of different perplexities is sufficient. If computing the t-SNE representations is the limiting factor, the perplexity can be calibrated for a subsample of the data, instead. [14] found that embedding a $\rho$-sample, where $\rho\in(0,1]$ is the sampling rate, with perplexity $\textrm{Perp}^{\prime}$ gives a visual impression of embedding the original data with perplexity
$\textrm{{Perp}}=\frac{\textrm{Perp}^{\prime}}{\rho}.$
With these concessions, applying this framework to calibrate perplexity is feasible for data sets of any size.
@@ 6 Discussion
We have illustrated the importance of acknowledging noise when performing dimension reduction by studying the roles perplexity and n_neighbors play in overfitting data. When using the original data to calibrate perplexity, our experiments agreed with perplexities previously recommended. When using just the signal, however, our experiments indicated that larger perplexities performed better. Low perplexities lead to overly-flexible t-SNE models that are heavily impacted by the presence of noise, while higher perplexities exhibit better performance due to increased stability. These considerations are especially important when working with heavily noised data, which are especially prevalent in the world of single-cell transcriptomics [5].
We have also presented a framework for modeling dimension reduction problems in the presence of noise. This framework can be used to study other hyperparameters and their relationships with noise. In the case when a specific signal structure is desired, this framework can be used to determine which dimension reduction method best preserves the desired structure. Further works should explore alternative methods for extracting the signal in way that preserves the desired structure.
@@ 7 Data Availability
Figure S1: Trefoil Plots
@@
Figure S2: Mammoth Plots
@@
SII.1 UMAP Plots (CyTOF)
Figure S3: UMAP Plots
SII.2 scRNA Dataset
This is a dataset of induced pluripotent stem cells generated from three different individuals [16]. The original data includes 864 units and 19,027 readings per unit. To process this zero-inflated count data, columns containing a large proportion of 0’s (20% or more) were removed before a log transformation was applied. This reduced the dimension to 5,431. A PCA pre-processing stop further reduced the dimension to 500, which still retained 88% of the variance of the log-transformed data. The signal was first taken to be the first five principal components, then the first 10 principal components. Notice the optimal perplexity when compared against the original data differed between these two experiments, even though it should theoretically be independent of the chosen signal dimension. This is due to the inherent randomness of the t-SNE algorithm.
Figure S4: scRNA Plots (r = 5)
Figure S5: scRNA Plots (r = 10)
SII.3 Microbiome Dataset
[2] compares the faecal microbial communities from 22 subjects using complete shotgun DNA sequencing. The original data contained 280 samples and 553 genera. To deal with a large number of near-zero readings, columns containing a large proportion of values less than $10^{-6}$ (60% or more) were removed. This reduced the dimension to 66. A PCA pre-processing was used to center and re-scale the data. The signal was first taken to be the first five principal components, then the first eight principal components. Notice the optimal perplexity when compared against the original data differed between these two experiments, even though it should theoretically be independent of the chosen signal dimension. This is due to the inherent randomness of the t-SNE algorithm.
Figure S6: Microbiome Plots (r = 5)
Figure S7: Microbiome Plots (r = 8)

### v5

@@ 4.2 Practical Examples
In addition to simulated data sets, we looked at three practical data sets: a single-cell RNA sequencing data set [20], a cytometry by time-of-flight (CyTOF) data set [21], and a microbiome data set [22]. For each data set, we compared the optimal perplexity for locally replicating the original data versus the estimated signal. We explore the scRNA-seq data set in detail here. The results of the other two practical examples can be found in Table 1. The details can be found in the Supporting Information (SI.1 and SI.2).
The scRNA-seq data set was generated from induced pluripotent stem cells collected from three different individuals. The original data includes 864 units and 19,027 readings per unit. To process this zero-inflated count data, columns containing a large proportion of 0’s (20% or more) were removed before a log transformation was applied. This step reduced the number of dimensions to 5,431. A PCA pre-processing step further reduced the number of dimensions to 500, which still retained 88% of the variance of the log-transformed data. Hence, the processed data set consisted of 864 observations in 500 dimensions, $Z+\epsilon\in\mathbb{R}^{864\times 500}$. To determine the dimensionality of the signal, we drew a scree plot (Figure 9). Note, the first eigenvalue (2359.357) was cut to fit the plot. A conservative estimate is five dimensions, so $Y$ was extracted by taking the first five principal components, $Y=\textrm{PCA}_{5}(Z+\epsilon)$. We computed the t-SNE representations for perplexities ranging from 10 to 280. For each perplexity, 20 different t-SNE representations were computed.
@@
As with the simulated examples, there is a difference in trend when switching the frame of reference (Figure 10). When compared against the original data, trustworthiness is maximized at a perplexity of 40, which is consistent with [4]’s recommendation of 5 to 50. When compared against the signal, trustworthiness is maximized at a larger perplexity of 120, reinforcing the hypothesis that lower values of perplexity may be overfitting the noise.
Visual inspection of the trustworthiness-maximizing representations reveals the effect of increasing the perplexity (Figure 11). A hierarchical clustering of the high-dimensional data was computed, then projected onto the trustworthiness-maximizing representations. Both representations depict a similar a structure, but the relative positioning of clusters differs. For example, the Class 1 cluster is the rightmost cluster in the perplexity = 40 representation, while the Class 5 cluster is the rightmost cluster in the perplexity = 120 representation. Furthermore, the left-to-right order of the Class 3, Class 8, and Class 9 clusters is reversed in both representations. In terms of local structure, the Class 3 and Class 7 clusters are better preserved in the perplexity = 40 representation, while the Class 12 cluster is better preserved in the perplexity = 120 representation. The perplexity = 40 representation also suggests the Class 2 cluster could potentially contain two separate clusters, but this is not consistent with the high-dimensional data according to the dendrogram and higher-order clusterings. The perplexity = 120 representation does not mislead in this way. Overall, the lower perplexity leads to tighter-knit clusters as expected. However, further investigation reveals the over-clustering may be unfaithful to the original data.
      Optimal Perplexity
      Data Set       $n$       $p$       $r$       signal + noise       signal
      Links [13]       500       10       3       40       80
      Trefoil [13]       500       10       3       35       100
      Mammoth [18]       500       10       3       30       80
      High-Dimensional Clusters       210       60       10       55       60
      scRNA-seq [20]       864       500       5       40       120
      scRNA-seq [20]       864       500       10       50       60
      CyTOF [21]       5,000       30       5       50       110
      CyTOF [21]       5,000       30       8       45       65
      Microbiome [22]       280       66       5       50       90
      Microbiome [22]       280       66       8       60       85
Table 1: Summary of Results
Figure 12: Trustworthiness vs. Perplexity for r = 10 (scRNA-seq)
Figure 13: Trustworthiness vs. n_neighbors for UMAP (scRNA-seq)
@@
If we, instead, decide to be more conservative and use the first 10 principal components to represent the signal, we still see a similar trend (Figure 12). Trustworthiness still increases then decreases with perplexity. When compared against the original data, trustworthiness is maximized at a perplexity of 50 (Note the optimal perplexity when compared against the original data differed between the two experiments, even though it should theoretically be independent of the chosen signal dimension. This is due to the inherent randomness of the t-SNE algorithm). When compared against the signal, trustworthiness is maximized at a perplexity of 60. By including three extra principal components in the signal, we’re assuming the data contains less noise, allowing the model to be more aggressive during the fitting process.
@@ 4.4 UMAP and n_neighbors
If n_neighbors functions similarly to perplexity, we’d expect small values of n_neighbors to overfit the data as well. An identical experiment was run using the Python package umap-learn [5] on the scRNA-seq data. n_neighbor values ranging from 10 to 300 were tested and an n_neighbors value of 190 maximized trustworthiness when comparing against the original data, but an n_neighbors value of 300 maximized trustworthiness when comparing against the signal (Figure 13).
@@ 5 Application
With a signal constructed, it remains to compute t-SNE/UMAP outputs at varying perplexities/n_neighbors. It’s recommended that at least a couple outputs are computed for each perplexity/n_neighbors to account for randomness in the algorithms. For each output, one must calculate the trustworthiness and Shepard goodness with respect to the signal. From there, one can choose the representation with the desired balance of local and global performance. A summary is given in Algorithm 1. Sample code is available at https://github.com/JustinMLin/DR-Framework/.
It is worth noting that computational barriers may arise, especially for very large data sets. To alleviate such issues, trustworthiness and Shepard goodness can be approximated by subsampling before calculation. Furthermore, t-SNE and UMAP are generally robust to small changes in perplexity and n_neighbors, so checking a handful of values is sufficient. If computing multiple low-dimensional representations is the limiting factor, one can try calibrating the hyperparameters for a subsample before extending to the full data set. [24] found that embedding a $\rho$-sample, where $\rho\in(0,1]$ is the sampling rate, with perplexity $\textrm{Perp}^{\prime}$ gives a visual impression of embedding the original data with perplexity $\textrm{{Perp}}=\frac{\textrm{Perp}^{\prime}}{\rho}$. With these concessions, applying this framework to calibrate hyperparameters should be feasible for data sets of any size.
6 Case Study
To demonstrate how one might apply this framework, we walk through a detailed case study on a modern scRNA-seq data set.
6.1 Data
Cryopreserved human peripheral blood mononuclear cells (PBMCs) from a healthy female donor aged 25 were obtained by 10x Genomics from AllCells. Granulocytes were removed by cell sorting, followed by nuclei isolation. Paired ATAC and Gene Expression libraries were generated from the isolated nuclei and sequenced. See [25] for details.
Figure 14: Scree Plot for PBMC Data Set
Figure 15: Trustworthiness vs. n_neighbors for UMAP (PBMC)
6.2 Pre-Processing
Pre-processing was completed using the $R$ package BPCells and the steps followed the provided tutorial [26] closely. Low quality cells (those that did not meet the required number of RNA reads, the required number of ATAC reads, or TSS Enrichment cutoffs) were filtered out before a matrix normalization was applied. The cleaned dataset contained 2,600 cells and 1,000 genes. The number of dimensions was then reduced to 500 using PCA, which retained 86% of the original variance. The processed data set to be analyzed contained 2,600 observations in 500 dimensions, $\mathbb{Z+\epsilon}\in\mathbb{R}^{2,600\times 500}$.
6.3 Determining the Signal
To determine the number of signal dimensions, a scree plot was drawn (Figure 14). The first eigenvalue was approximately 188 but was trimmed to fit the plot. Four dimensions, a relatively conservative estimate, were chosen to represent the signal, $Y\in\mathbb{R}^{2,600\times 4}$.
Figure 16: Cell Types (PBMC)
Figure 17: Plot of Dendritic Cells (PBMC)
6.4 Results
UMAP was applied with multiple values of n_neighbors. 20 representations were computed for each value, and trustworthiness was measured with respect to both the entire data and the signal. Trustworthiness was maximized at a n_neighbors value of 50 when comparing against the entire data and a value of 70 when comparing against the signal (Figure 15). Cell types (B, T, Monocyte, NK, Dendritic cell, CD8 T) were assigned to each cluster by exploring marker genes (Figure 16). See [26] for details.
Figure 18: PCA Applied to Dendritic Cells (PBMC)
Figure 19: PCA applied to Dendritic Cells (PBMC)
6.5 Analysis
In all three representations, the primary division of cell types is between monocytes and some of the dendritic cells vs. the T, B, CD8 T, and NK cells. In the default UMAP representation, the B cells form a cluster that is quite distinct from the T, CD8 T, and NK cells. As we increase n_neighbors to 50 and 70, the B cell cluster moves closer to the T/CD8 T/NK cell cluster. The closer proximity of the B cells to the T, CD8 T, and NK cells in the n_neighbors = 70 representation is consistent with the over-arching categorization of T, CD8 T, NK, and B cells as lymphocytes, as opposed to monocytes.
Perhaps a starker difference between the representations concerns the dendritic cells (DCs). In the n_neighbors = 15 and n_neighbors = 50 representations, there are three distinct clusters of DCs, whereas there are only two in the n_neighbors = 70 representation (Figure 17). Principal component analysis of the DCs alone suggests that the DCs are either two clusters, one of which is more diffuse than the other, or three clusters, two of which are fairly close together (Figure 18). Standard metrics for determining the number of clusters suggest the same. The silhouette width metric suggests two clusters, while the gap statistic suggests three (Supporting Information SII.3). However, the three-cluster solution given by k-means and visual inspection of the principal components plot does not align with the three clusters in the n_neighbors = 15 or n_neighbors = 50 representation. The green and orange clusters are represented faithfully, but the third, more diffuse, purple cluster is split across two DC clusters in the n_neighbors = 15 and n_neighbors = 50 representations (Figure 19). The degree of separation is much lesser in the n_neighbors = 70 representation. Therefore, the n_neighbors = 15 and n_neighbors = 50 representations inaccurately represent the dendritic cells in a way that the n_neighbors = 70 representation does not.
@@ 7 Discussion
We have illustrated the importance of acknowledging noise when performing dimension reduction by studying the roles perplexity and n_neighbors play in overfitting data. When using the original data to calibrate perplexity, our experiments agreed with perplexities previously recommended. When using the signal, however, our experiments indicated that larger perplexities perform better. Low perplexities/n_neighbors lead to overly-flexible models that are heavily impacted by the presence of noise, while higher perplexities/n_neighbors exhibit better performance due to increased stability. These considerations are especially important when working with heavily noised data, which are especially prevalent in the world of single-cell transcriptomics [27].
We have also presented a framework for modeling dimension reduction problems in the presence of noise. This framework can be used to study other hyperparameters and their relationships with noise. In the case when a specific signal structure is desired, this framework can be used to determine which dimension reduction method best preserves the desired structure. Further works should explore alternative methods for extracting the signal a in way that preserves the desired structure.
@@ 8 Data Availability
For the trefoil example, the signal $Y$ consisted of a trefoil knot embedded in three dimensions containing 500 points. $Z+\epsilon$ was constructed by adding seven superfluous dimensions and isotropic Gaussian noise. Various degrees of noise were tested ($sd=5,10,15,20,25,30$). The first two plots depict Trustworthiness vs. Perplexity and the trustworthiness-maximizing embeddings for the $sd=10$ case. The third plot shows the trustworthiness-maximizing perplexity for the different degrees of noise.
Figure S1: Trustworthiness vs. Perplexity (Trefoil)
Figure S2: Trustworthiness-Maximizing Representations (Trefoil)
Figure S3: Optimal Perplexity (Trefoil)
@@
For the mammoth example, the signal $Y$ consisted of 500 points in three dimensions. The data was randomly sampled from the mammoth data set used in [18]. $Z+\epsilon$ was constructed by adding seven superfluous dimensions and isotropic Gaussian noise. Various degrees of noise were tested ($sd=0.5,1,1.5,2,2.5,3$). The first two plots depict Trustworthiness vs. Perplexity and the trustworthiness-maximizing embeddings for the $sd=1$ case. The third plot shows the trustworthiness-maximizing perplexity for the different degrees of noise.
Figure S4: Trustworthiness vs. Perplexity (Mammoth)
Figure S5: Trustworthiness-Maximizing Representations (Mammoth)
Figure S6: Optimal Perplexity (Mammoth)
@@
SII.1 CyTOF Data Set
The CyTOF data set contained 239,933 observations in 49 dimensions [21]. To reduce the computational load, a subset of 5,000 observations was sampled. A log transformation was followed by a PCA pre-processing step to reduce the number of dimensions to 30, which still retained 77% of the variance in the original data. The processed data set to be studied consisted of 5,000 observations in 30 dimensions. The signal was first taken to be the first five principal components, then the first eight principal components. A hierarchical clustering of the high-dimensional data was computed then projected onto the trustworthiness-maximizing representations.
Figure S7: Trustworthiness vs. Perplexity for r = 5 (CyTOF)
Figure S8: Trustworthiness-Maximizing Representations for r = 5 (CyTOF)
Figure S9: Trustworthiness vs. Perplexity for r = 8 (CyTOF)
Figure S10: Trustworthiness-Maximizing Representations for r = 8 (CyTOF)
SII.2 Microbiome Data Set
[22] compares the faecal microbial communities from 22 subjects using complete shotgun DNA sequencing. The original data contained 280 samples and 553 genera. To deal with a large number of near-zero readings, columns containing a large proportion of values less than $10^{-6}$ (60% or more) were removed. This reduced the dimension to 66. A PCA pre-processing was used to center and re-scale the data. The signal was first taken to be the first five principal components, then the first eight principal components. A k-means clustering of the high-dimensional data was computed then projected onto the trustworthiness-maximizing representations.
Figure S11: Trustworthiness vs. Perplexity for r = 5 (Microbiome)
Figure S12: Trustworthiness-Maximizing Representations for r = 5 (Microbiome)
Figure S13: Trustworthiness vs. Perplexity for r = 8 (Microbiome)
Figure S14: Trustworthiness-Maximizing Representations for r = 8 (Microbiome)
SIII PBMC Data Set
Figure S15: Average Silhouette Width for Dendritic Cells
Figure S16: Gap Statistic for Dendritic Cells